Calculate the mole of for chemicals below.

155g of Ca3(PQ4)z

I0g of chlorine gas

0.5g of Sodium hydroxide



1
Expert's answer
2021-11-25T01:30:11-0500

Molar mass of Ca3(PQ4)z= 310.18

= 155/310.18=0.4997mol

Molar mass of chlorine gas= 70.9

10/70.9=0.141mol

Molar mass of NaOH =39.997

= 0.5/39.997=0.0125mol